OpenStudy (anonymous):

simpifly -1 3/7 x (-3 2/3)

OpenStudy (anonymous):

use parenthesis.

OpenStudy (anonymous):

\[-1\frac{ 3 }{ 7 }\times-3\frac{ 2 }{ 3 }??\] Simplify the fractions by finding common denominators \[-(\frac{ 7 }{ 7 }+\frac{ 3 }{ 7 })\times-(\frac{ 9 }{ 3 }+\frac{ 2 }{ 3 })\] \[\frac{ -10 }{ 7 }\times \frac{ -11 }{ 3 }\] Multiply the numerators and denominators \[\frac{ 110 }{ 21 }\]